### v3

@@ 0.2.5 Sex classification
0.2.5.1   Forced site-target independence
@@ 0.2.6 Dementia and mild cognitive impairment classification
0.2.6.1   Forced site-target dependence
@@
In our experiments using the ADNI dataset, where the data were acquired in different sites, we selected 100 dementia-MCI patients and 10 controls from one site. We selected 100 control patients and 10 dementia-MCi patients from another site, again forcing the site-target relationship (Table 10). The images were processed with FreeSurfer 46. The thickness from 74 cerebral and sub-cerebral structures were extracted as features.
@@
0.2.6.2   Forced site-target independence
Using the ADNI dataset the same extracted features were used. From the dataset, 126 dementia-MCI and control patients were randomly selected from the first site while 56 dementia-MCI and control patients were randomly selected from the second site (Table 11).
@@ 0.2.8 Forced site-target dependence
The eICU 32, 31, 47 dataset was used for the experiments, which contains 200859 ICU stays from 139367 patients in 208 different ICUs across the United States. We use a well-known problem of classified hospital discharge (Expired or Alive), in septic patients 48, 49. The approach described in 39 was followed for selecting the features and extracting the patient cohort. The features used were arterial blood gases: paO2, paCO2, pH, base excess, Hgb, glucose, bicarbonate, and lactate. After the patients’ selection, a final dataset of 496 Expired and 3021 Alive patients was retained. From this filtered dataset, we remove those sites with less than 50 stays, retaining 20 final sites.
From 20 of these sites, all “Alive” patients, except for one, were removed for the 10 sites with more “Expired” patients. Contrary, all the “Expired” patients, except for one, were removed from the 10 sites with less number of “Expired” patients. Note that in this case, the classes (Alive and Expired) are represented in several sites and are not the same in each site (Table 12), compared with the previous classification experiments performed. A total of 249 Expired and 666 Alive patients were used in this experiment.
@@ 0.2.9 Forced site-target independence
From the eICU, the same features extraction and patient selection were made (496 Expired and 3021 Alive patients). The same 20 sites, with more than 50 images, were used.
@@ Methods
The proposed method rests on the use of “pretended” target values on the harmonization process, thereby enabling predictions without requiring test target values (Figure 2). For our experiments, the available data is divided into train and test, simulating a real use case. The training fold is further divided into inner train and validation folds. A neuroHarmonize model is trained on the inner training data to learn to remove the site’s effect. The inner training data is harmonized and a Predictive model is trained on the harmonized inner train data to predict the target. This model needs to be chosen as the best possible model to solve the problem at hand.
Using the trained neuroHarmonize model, the validation samples are harmonized while “pretending” their target value. For example, for a binary classification problem, all the validation labels are set as the first class, pretending that all validation samples belong to the first class. Using these “pretended” labels, the validation data is harmonized and a prediction is made using the trained Predictive model. Later, the validation labels are set to the second class and the validation data is harmonized again and a new prediction is generated. In general, for a classification task, the set of available classes is pretended, while for a regression task, the values are linearly sampled in the target’s range. All the predictions, generated with the pretended harmonized data, are concatenated and a “Score matrix” is created. This matrix has a dimension of number of validation samples times the number of labels. To effectively utilize the training dataset, a K-fold cross-validation (CV) procedure was employed, generating out-of-sample predictions for the entire dataset. Using this Score matrix as input features, a “Stack” model is trained to predict the target and give a final prediction.
At the test time, when the test label is not available, the same procedure is followed. For example, in a binary classification problem, a test sample will be harmonized using the neuroHarmonize model first pretending that the test label belongs to the first class. The Predict model will generate a prediction using the harmonized data and the process will be repeated pretending the test sample belongs to the second class. Both predictions generated by the Predictive model are concatenated and a test Score matrix is built. This matrix is used by the Stack model, which generates the final prediction.
@@ 0.3 Machine learning model
For the binary classification problem using the synthetic data (MAREoS datasets), a Random Forest model (RF) 50, with default sklearn parameters, was used as a Predictive Model, and a Logistic Regression (LG) 51 was used as a Stack Model. The same RF model was used to train a model with the original data to obtain a classification baseline for each dataset (Baseline model).
For the age regression problems using real MRI data, Relevance Vector Regression 52 with a polynomial kernel of degree 1 (RVR) was used as a Predictive and Stack model for PrettYharmonize. The RVR model was used for the rest of the harmonization schemes 36. A 5-fold cross-validation scheme was used, and the Mean Absolute Error (MAE), coefficient of determination (R2), and age bias (Pearson’s correlation between the true age and the difference between the predicted and true age) were calculated on the test sets.
For the sex classification using real MRI data, RVR was used as a Predictive and Stack model. A 5 times repeated 5-fold stratified cross-validation scheme was used, and the Area under the receive operation curve (AUC), balanced accuracy (bACC) and were calculated on the test

### v4

@@ 0.2.5 Sex classification
Forced site-target independence
@@ 0.2.6 Dementia and mild cognitive impairment classification
Forced site-target dependence
@@
In our experiments using the ADNI dataset, where the data were acquired in different sites, we selected 100 dementia-MCI patients and 10 controls from one site. We selected 100 control patients and 10 dementia-MCi patients from another site, again forcing the site-target relationship (Table 10). The images were processed with FreeSurfer [46]. The thickness of 74 cerebral and sub-cerebral structures were extracted as features.
@@
Forced site-target independence
Using the ADNI dataset, the same extracted features were used. From the dataset, 126 dementia-MCI and control patients were randomly selected from the first site while 56 dementia-MCI and control patients were randomly selected from the second site (Table 11).
@@ 0.2.8 Forced site-target dependence
The eICU [47, 48, 49] dataset was used for the experiments, which contains 200859 ICU stays from 139367 patients in 208 different ICUs across the United States. We use a well-known problem of classified hospital discharge (Expired or Alive), in septic patients [56, 57]. The approach described in [32] was followed for selecting the features and extracting the patient cohort. The features used were arterial blood gases: paO2, paCO2, pH, base excess, Hgb, glucose, bicarbonate, and lactate. After the patients’ selection, a final dataset of 496 Expired and 3021 Alive patients was retained. From this filtered dataset, we remove those sites with less than 50 stays, retaining 20 final sites.
From 20 of these sites, all “Alive” patients, except for one, were removed for the 10 sites with more “Expired” patients. On the contrary, all the “Expired” patients, except for one, were removed from the 10 sites with fewer “Expired” patients. Note that in this case, the classes (Alive and Expired) are represented in several sites and are not the same in each site (Table 12), compared with the previous classification experiments performed. A total of 249 Expired and 666 Alive patients were used in this experiment.
@@ 0.2.9 Forced site-target independence
The same feature extraction and patient selection were made (496 Expired and 3021 Alive patients) from the eICU. The same 20 sites, with more than 50 images, were used.
@@ Methods
The proposed method introduces a new harmonization scheme based on the neuroHarmonize model. This scheme relies on the use of “pretended” target values during the harmonization process, enabling predictions to be made without requiring actual test target values (Figure 2). For our experiments, the available data is split into training and test sets, simulating a real-world use case. The training fold is further divided into inner training and validation folds. A neuroHarmonize model is trained on the inner training data to learn how to remove the site’s effect. The inner training data is then harmonized, and a predictive model is trained on the harmonized inner training data to predict the target. This predictive model must be selected as the best possible model to address the specific problem at hand.
Using the trained neuroHarmonize model, the validation samples are harmonized while “pretending” their target value. For example, for a binary classification problem, all the validation labels are set as the first class, pretending that all validation samples belong to the first class. Using these “pretended” labels, the validation data is harmonized and the predictions are made using the trained Predictive model. Later, the validation labels are set to the second class and the validation data is harmonized again and a new prediction is generated. In general, for a classification task, the set of available classes is pretended, while for a regression task, the values are linearly sampled in the target’s range. All the predictions, generated with the pretended harmonized data, are concatenated and a “Score matrix” is created. This matrix has a dimension of number of validation samples times the number of labels. To effectively utilize the training dataset, a K-fold cross-validation (CV) procedure was employed, generating out-of-sample predictions for the entire dataset. Using this Score matrix as input features, a “Stack” model is trained to predict the target and give a final prediction.
At the test time, when the test label is not available, the same procedure is followed. For example, in a binary classification problem, a test sample will be harmonized using the neuroHarmonize model first pretending that the test label belongs to the first class. The Predict model will generate a prediction using the harmonized data, and the process will be repeated pretending the test sample belongs to the second class. Both predictions generated by the Predictive model are concatenated and a test Score matrix is built. This matrix is used by the Stack model, which generates the final prediction.
@@ 0.3 Machine learning model
For the binary classification problem using the synthetic data (MAREoS datasets), a Random Forest model (RF) [40], with default sklearn parameters, was used as a Predictive Model, and a Logistic Regression (LG) [39] was used as a Stack Model. The same RF model was used to train a model with the original data to obtain a classification baseline for each dataset (Baseline model).
For the age regression problems using real MRI data, Relevance Vector Regression [38] with a polynomial kernel of degree 1 (RVR) was used as a Predictive and Stack model for PrettYharmonize. The RVR model was used for the rest of the harmonization schemes [31]. A 5-fold cross-validation scheme was used, and the Mean Absolute Error (MAE), coefficient of determination (R2), and age bias (Pearson’s correlation between the true age and the difference between the predicted and true age) were calculated on the test sets.
For the sex classification using real MRI data, RVR was used as a Predictive and Stack model. A 5 times repeated 5-fold stratified cross-validation scheme was used, and the Area under the receiver operating curve (AUC), balanced accuracy (bACC) and were calculated on the test